OpenStudy (anonymous):

Solve the equation or formula for the indicated variable. R = ts^2 + 3 for s

OpenStudy (anonymous):

I think I know how to solve it already and the answer I just am hoping to verify my theory.

OpenStudy (anonymous):

I believe the answer is s=(under radical)R-3/t. The t is not under the radical

OpenStudy (anonymous):

\[\frac{ R }{ t } = s ^{2} + 3\] \[\frac{ R }{ t } - 3 = s ^{2} \] \[\sqrt{\frac{ R }{ t } - 3} = s\] This is what I've come up with

OpenStudy (anonymous):

Alright how did you come to this conclusion?

OpenStudy (anonymous):

Turns out my answer was correct.